CRAZYkillerboy:

How to do this prob step by step? (1 + -2) x2 + (2+3)x + (1+4)

steevh:

(2+(-4) + 2x+3x + 5=0 (2-4)+5x+5=0 -2+5x+5=0 5x+5-2=0 5x+3=0 5x=-3 x=-3/5

CRAZYkillerboy:

is that right?

steevh:

yup

CRAZYkillerboy:

ooo thx

steevh:

easy

CRAZYkillerboy:

hard!!

CRAZYkillerboy:

thx bro

steevh:

yw